# VSF: Simple, Efficient, and Effective Negative Guidance in Few-Step Image Generation Models By Value Sign Flip

### v2

@@ 4.11 Computational Cost
From the table, VSF requires marginally more time and memory than NAG in SD3.5, while they are both significantly faster than theoretical CFG time, which would be twice the baseline. In Wan, VSF outperforms NAG and is even slightly better than the baseline (likely due to nature variation) in terms of compute time, though it consumes 1GB more memory, likely due to the attention bias being stored. Since this bias is optional, we tested VSF Wan’s performance with it removed, which results in an improvement in VRAM usage such that it uses the same amount of VRAM as baseline and NAG, and no change in runtime.
5 Conclusion and Future Work
In this paper, we introduced VSF, a novel approach for enhancing negative prompt adherence in image and video generation models. Our method involves flipping the sign of attention values corresponding to negative prompts, effectively suppressing unwanted content. Experimental results indicate that VSF significantly outperforms previous methods, NAG Chen et al. (2025a), in terms of negative prompt adherence, with only minor trade-offs in overall quality and positive prompt fidelity. VSF also only has one main hyperparameter and one minor hyperparameter, making it easier to tune them in downstream tasks.
@@ 5 Conclusion and Future Work
Future work may involve extending VSF to other architectures, such as Flex Schnell, improving robustness through normalization and blending techniques similar to those employed by NAG, and optimizing computational efficiency by using a better attention implementation. Additionally, we observed some inaccuracies in MLLM judgment due to ambiguities or minimal differences in visual differences. Conducting a larger-scale human evaluation study would help mitigate inaccuracies observed in MLLM-based assessments. Investigating the attention maps and diffusion trajectories of our model could further elucidate the underlying mechanisms of VSF.
Acknowledgements
Appendix A Appendix
A.1 Qualitative Results for Wan
To qualitatively evaluate the effects of our method, we present several example generations across different models. These samples are intended for visualization, and results may vary with different hyperparameters or sampling conditions. The examples demonstrate both strengths and limitations across models and are not meant to reflect comprehensive performance.
In the first example, we evaluate cutting tomatoes on a board, where the negative prompt is a wooden board. The original video (generated without negative guidance) includes a wooden board. However, the NAG and VSF outputs both replace it with non-wooden surfaces—a plastic or glass-like surface. However, both original and NAG outputs exhibit unnatural physics during the cutting motion. The VSF sample avoids major unnatural physics but fails to preserve the shape of the tomato, as the tomato deforms unrealistically during slicing. These issues likely arise from limitations in the small 1.3B parameter model since it also appears in the original video.
The second example involves a lava river without glowing. In still frames, the outputs from both NAG and VSF appear similar to regular rivers, but in motion, they exhibit texture and flow characteristics resembling lava. The NAG sample, however, leans more toward a natural river, reflecting weaker adherence to the intended appearance.
Figure 11: Example of Wan 2.1 with no negative guidance (NONE), NAG guidance (NAG), and our guidance (VSF). Positive and negative prompts are shown at the bottom of each figure.
In the third example, the prompt requests a plane with no wings. The original outputs clearly retain wings. The NAG output reduces the wing size but does not remove it completely. VSF most closely satisfies the constraint, with wings nearly absent, indicating a stronger capability in element removal.
The fourth example aims to remove bias in generating a machine learning scientist by excluding the common depiction of a male with glasses. Both NAG and VSF successfully eliminate the targeted attributes. Among them, NAG produces the more natural-looking output, with better visual coherence.
In the fifth case, the prompt asks for a pet running through flowers, explicitly excluding dogs. The original sample resembles a dog-like hybrid. The NAG version trends more cat-like and adheres better to the constraint. The VSF sample avoids the dog but results in a character resembling a fictional figure from a children’s story, making its alignment with the “pet” concept weaker.
In the sixth example, we request a cat chef cooking in a kitchen without a window. Since the original sample already lacks a window, it does not test model behavior under the constraint. In the NAG output, under such a high guidance scale, it introduces undesired concepts, such as a human inside the cat adding seasoning to the pan, suggesting instability under strong steering.
The final example prompts for a laptop playing a cat video without a keyboard. Both NAG and VSF fail to remove the keyboard, likely due to its strong association with the laptop concept. Additionally, the NAG output results a cat with two heads in the screen, which may stem from the high guidance scale or the model’s limited capacity.
From these examples, we observe that both models succeed and fail in different scenarios. VSF is more effective at removing explicit elements (e.g., wings) but less reliable at excluding abstract sub-concepts (e.g., dog from pet, male from scientist). Overall, the outputs of VSF and NAG are comparable. However, our method (VSF) operates at a lower compute cost, as discussed in the following section.

### v5

@@ Appendix H Computational Cost
From the table, VSF requires marginally more time and memory than NAG in SD3.5, while they are both significantly faster than theoretical CFG time, which would be twice the baseline. In Wan, VSF outperforms NAG and is even slightly better than the baseline (likely due to nature variation or noise) in terms of compute time, though it consumes 1GB more memory, likely due to the attention bias being stored. Since this bias is optional, we tested VSF Wan’s performance with it removed, which results in an improvement in VRAM usage such that it uses the same amount of VRAM as baseline and NAG, and no change in runtime.
Appendix I External Baselines
In addition to other guidance methods applied to SD-3.5-turbo, we evaluated several external baselines. The first baseline employs a generate-then-edit approach, loosely inspired by Generate-Plan-Edit (GraPE) (Goswami et al., 2025), but omitting the planning stage as our goal is straightforward (removing unwanted elements). Specifically, we first generated images using SD-3.5-Large-Turbo without a negative prompt, and subsequently edited out the unwanted elements automatically using Flux Knoest (Labs et al., 2025), an image editing model, using prompt Remove [negative prompt].
The second baseline utilizes GPT-4o’s native image generation capability. GPT-4o has demonstrated strong prompt-following performance(Wei et al., 2025), including in handling negation tasks. As GPT-4o lacks explicit negative prompt functionality, we formatted prompts as [Positive prompt], but with no [negative prompt]. Since our focus is on evaluating negation rather than image quality, we adopted the “low” generation setting. Besides GPT-4o, we also added the newly released Nano Banana from Google. It is also a language model-based image generation model and has received a good reputation in the image generation community.
The third baseline we included is Janus-4o (Chen et al., 2025b), a model distilled from GPT-4o onto the Janus-Pro base architecture (Chen et al., 2025c). Given GPT-4o’s strong prompt-following performance, we anticipated competitive results from Janus-4o. We provided negative prompts directly as negations within the positive prompts, same as GPT-4o.
All measure time is measured on Google Colab 40GB A100 GPU, and for Qwen-Image and Generate+Edit, model CPU offloading is enabled.
The results are presented in Table 1 in the main text. The table indicates that VSF Strong achieves the second-best negative score, only behind GPT-4o, while also demonstrating a significantly faster runtime compared to all other methods, outperforming even the generate-then-edit pipeline. The GPT-4o distilled model, Janus-4o, has an unexpectedly low negative score, which could be because they did not have enough negation-included prompts in the distillation data. The VSF Quality had a lower negative score compared to Generate+Edit, while having a much higher positive and quality score, and shorter runtime.
Appendix J SD3.5-Large-Turbo Qualitative Results
Figure 12: Selected Results for Comparison. Positive promtes are condensed for spacing.
Selected qualitative results are shown in Figure 12. The positive prompt is condensed for spacing. For the glasses without lens images, both NAG and NAG Strong generated classes clearly have a lens. For the VSF-generated image, we can see the lens is missing, even though the frames are floating. However, this issue was also presented in NAG Strong’s image, even though it still has glasses. For a sailboat without sails, all other methods generated smaller but still existing sails, while VSF successfully avoided sails. In the third image of a lighthouse without a lamp, both VSF and NASA have no visible lamp, while the images from NAG and NAG Strong have a clear lit lamp. In the image of a bicycle without a chain, NASA generated a blurry image without bikes at all, while NAG generated a normal image, and NAG Strong generated a slightly distorted image of a bike with no seat yet the chain is still present. VSF successfully generated a bike without a chain, even though it also removed the seat. For the prompt of a lantern with no glass panes, NASA generated a lamp with frosted panes, NASA++ generated a classic glass pane, and NASA++ generated broken frosted panes. VSF, in this case, generated a pane that is clearly not glass. In the last example of a T-shirt, NASA generated a blurry image with still one sleeve visible, and NAG generated the T-shirt with both sleeves visible. NAG Strong and VSF both avoided the sleeve, even though NAG Strong has some artifacts.
Appendix K Qualitative Results for Wan
In Figure 13, we showed 3 examples generated from Wan-2.1-14B. In the first example, we successfully removed the stars in the background while keeping other elements intact. On the right side, in the absence of stars, the moon lander is generated to fill the space. In the second video, we successfully generated a windowsill without a curtain. In the last video, the generated video from VSF contains no trees on the left, and instead, it fills it with a hill. There are still some bushes on the right side, which do not violate the negative prompt of “trees. All videos have the same high quality as the original one.
Figure 13: Qualitative Results for Wan
Appendix L Failure Cases
Figure 14: Failed examples, positive prompts are condensed for spacing.
Like any method, our method is not perfect, especially in a challenge dataset like NegGenBench. In Figure 14, we showed 4 failed cases. In the case where we want to generate a keyboard without a spacebar, the generated object is technically a key-board (an array of keys) and has no space bar, but it is not what people imagine when they think about ”keyboard with no spacebar.” The second failed case is another image from glasses with no lens; the generated image has no lens, but the frame is twisted in an unnatural way. In the third example, where a house with no roof is needed, VSF completely missed the negative prompt, possibly due to the strong association between roof and house. In the last example of a cat with no whiskers, the generated image technically has no whiskers and looks like a cat, but it looks more like a cat statue instead of a living cat.
Appendix M Non-Object Negative Guidance
In this paper, we focused on removing a critical component in the image. To further validate our negative guidance method in other areas, we also tested it on style avoidance. In Figure 5 (in main text), we show four examples, each of which is generated using the same seed. We can see that when prompted with famous artwork (e.g.,”A painting of Starry Night from the 1890s” or ”Mona Lisa oil painting”) but with a negative prompt of the artist’s name style, the generated image avoided the specific art style but kept the semantic meaning of the positive prompt. When prompted to give an old photo but not monochrome, the generated image is more like an old-style color photo, follows both non-monochrome and also not very bright (as old photos, even in color, are less vibrant). We find these examples interesting and think they can be used for machine unlearning, using a similar method as in (Gandikota et al., 2023).
Appendix N An Experiment on Anti-Aesthetics Arts
Current image generation models are typically finetuned to align with so-called “human preference.” However, we argue that there is no universal standard for human preference, and it cannot be defined solely by developers, who inevitably bring their own interests and assumptions. Aligning models exclusively with such values risks introducing bias and potentially marginalizing minority perspectives and interests (Arzberger et al., 2024; Turchin, 2019; Sutrop, 2020; Guo et al., 2025).
In the context of image generation, this alignment may lead to homogenization of style or taste, producing only broadly pleasing outputs for the general population. Such uniformity can suppress niche demands for degraded, low-quality, or unconventional aesthetics. To counteract this, one possible approach is the use of negative guidance to steer outputs away from mainstream preferences. In this experiment, we tested how VSF can address this issue. We ran our VSF in settings where $\alpha=0$, which shows on the left, and $\alpha\in[0,4]$, which shows on the right side. The image with $\alpha=0$ might not be the same as the one without guidance, but should be an image without negative guidance.
Figure 15: Image with abstract style receives a lower score in HPSv2 (reflecting human preference; traditionally, models aim for higher scores).
The first test used prompts containing the same object in both positive and negative form, with the goal of producing abstract art. This works by semi-canceling the main object, making it appear in an abstract form. Abstract styles are often disadvantaged in alignment settings, since reward models typically favor realistic or figurative outputs. VisionReward (Xu et al., 2025) encodes this bias through its scoring metric, and LAPIS (Maerten et al., 2025) reports that abstract paintings generally receive lower preference scores. Figure 15 shows that an abstract image gets a much lower score compared with a figurative one. As shown in the first two rows of Figure 18, the apple, people, and cat appear in abstract form, demonstrating a clear shift away from the default figurative tendency of the aligned models when VSF is applied. For the last image of a dog, we used “cute” as a negative prompt, which usually describes realistic objects, and we achieved a very abstract and artistic image. Figure 16 shows how abstract the image gets as the scale increases. More examples are shown in Figure 17.
Figure 16: Abstraction of the image as scale increases.
Figure 17: More Abstract Arts Examples
In the second test, the goal was to diverge from styles that are generally appreciated. The positive prompt specified the desired style, while the negative prompt contained descriptions of commonly preferred styles. Importantly, the positive prompt clearly described the intended output, so a faithful model should follow it rather than default to generalized human preference. The tested cases included desaturated color, sad emotion, pixelated art, insufficient lighting, unnatural colors, and a non-beautiful cat. Results show that the baseline model struggled to maintain these characteristics, often reverting to conventionally “beautiful” imagery, whereas VSF successfully produced outputs aligned with the specified unconventional styles.
Figure 18: A test for anti-aesthetics. The left image is generated with $\alpha=0$, and the right image is generated with $\alpha>0$. These tests aim to move away from universally pleasing styles and demonstrate the ability to capture more diverse aesthetic preferences.
Appendix O Negation-Aware MLLM
Upon visual inspection, we observed that GPT-4o effectively avoids many negative prompts, though occasionally the ambiguity within negative prompts (e.g., the term “door” referring either to the door panel or the entrance) and the vision ability of MLLM itself (i.e., hallucination) leads to lower negative scores. It is possible that the negative scores for other methods might also be underestimated. We acknowledge this as a limitation associated with using an MLLM as the evaluator. Thus, we provided a better negotiation understanding of MLLM and used this MLLM to evaluate different guidance methods.
To enable future research on evaluating negation prompting in generative models, a reliable and fast (LLaMA Maverick is too big) evaluation model is necessary. Previous CLIP-based studies concentrated on simple negations (e.g., ”a cat that is not on the grass”) rather than more complex cases such as those in NegGenBench. To address this, we finetuned a multimodal large language model, Qwen-2.5-VL-7B, called NegAwareQwen for improved understanding.
We created 100 additional prompts using GPT-5, each paired with a positive and a negative pair and 2 questions. For each prompt, we generated two images with the three models (NAG, VSF Quality, and NASA) and selected 722 for manual scoring on two dimensions: adherence to the negative prompt and overall quality. We did not assess the positive prompt evaluation as those are simpler, and almost all images in the dataset have a perfect positive prompt following. Negative adherence was rated on a three-level scale: 0 (ignored), 0.5 (partial), and 1 (fully followed). When generating the samples, we slightly randomly adjust the hyperparameter in a small range to create more diverse data (For VSF, $\alpha=3.3\pm 1,\beta=0.2$; for NASA, $\alpha=0.15\pm 0.05$; for NAG, $\phi=8\pm 4,\alpha=0.5\pm 0.2,\tau=4\pm 2$). Since this makes image generation models generate sub-optimal images, the rating results of each model’s images are not used for direct comparison. The dataset and the model will be opened after publication.
Note that here Llama showed a very weak r-score for quality; this is because all the images are evaluated using relatively high-quality images (unlike in the ablation study, where many images are lower quality). We did not compare the 7B untrained model because it often failed to output the structure data needed.
Table 7: Negation-Aware LLM Evaluation on Testing Set
Parameters r ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$)
Llama Maverick 400B 0.05 0.83 0.59
Llama Maverick CoT 400B 0.03 0.77 0.50
Qwen-2.5-VL 32B 32B 0.28 0.80 0.36
Qwen-2.5-VL 32B CoT 32B 0.31 0.88 0.70
NegAwareQwen-7B 7B 0.37 0.86 0.65
NegAwareQwen-32B 32B 0.34 0.90 0.76
The model was finetuned using prompts from the dataset of all 3 models. We trained the model using QLoRA Dettmers et al. (2023) with rank of 8 and $r=8$, applied to query, key, value projections in both the vision encoder and language model with dropout of 0.1. Model is trained using $lr=5\times 10^{-5}$ (with warm up and decay), WeightDecay=0.1, BatchSize=16, Epoch=5. The dataset is split into train-val with a 90-10 ratio based on the prompt level splitting and aiming for balanced scores in each split. We treat 0.5 as False and calculate negative scores as a binary metric. Results are presented in Table 7 with comparison with the same model without finetuning and LLaMA-Maverick.
Appendix P Evaluation Using NegAwareQwen
We re-evaluated VSF Quality, NAG, NAG Strong, NASA, GPT-4o, and Nano Banana images generated using the prompts and seeds in the main text of the paper using our finetuned NegAwareQwen; the results are shown in Table 8. We did not round the 0.5 score. We used a positive score from the original Table 2 as our finetuned version was trained on largely positive compliance samples. The results match our observation with human validation and MLLM evaluation, that our method gets the highest negative score while having better or comparable positive and quality scores with other open source guidance methods.
Table 8: Evaluation of Different Methods using Our NegAwareQwen-32B
Positive Score ($\uparrow$) Negative Score ($\uparrow$) Quality Score ($\uparrow$)
VSF Quality 0.980 0.330 0.814
VSF Strong 0.870 0.415 0.814
NASA 0.950 0.224 0.727
NAG Strong 0.950 0.168 0.795
NAG 1.000 0.147 0.812
GPT-4o 0.978 0.619 0.812
Nano Banana 0.985 0.406 0.817
Appendix Q Abliation Study On $\alpha$ and $\beta$
To study the effects of $\alpha$ and $\beta$ and hyperparameter sensitivity, we studied the effects of the two hyperparameters. We used 30 randomly selected prompts from the dataset and tested the effects of $\alpha$ and $\beta$ on the positive, negative, and quality scores. When testing the effects of $\alpha$, we set $\beta$ to 0, and when studying the effects of $\beta$, we set $\alpha$ to 3.5. All generations use the same seed. The images are evaluated using our NegAwareQwen-32B running for negative and quality score and used LLaMA for the positive score. Qualitative results of the effects of $\alpha$ and $\beta$ are shown in Figure 10 and quantitative results are shown in Figure 19. We can see that as $\alpha$ increases, negative scores increase while positive scores decrease. When $\beta$ increases, the negative score decreases while the positive score increases, which could be noise. In both cases, the quality scores only change slightly.
Figure 19: Relationships between metrics and $\alpha$ and $\beta$
Appendix R Future Work
@@ Appendix I External Baselines
Future work may involve applying it to non-diffusion models (like Janus-4o (Chen et al., 2025b)) or models with complex text encoders (like Qwen-Image (Wu et al., 2025)), improving robustness through normalization and blending techniques similar to those employed by NAG, and optimizing computational efficiency by using a better attention implementation. Additionally, we observed some inaccuracies in MLLM judgment due to ambiguities or minimal differences in visual differences. Conducting a larger-scale human evaluation study would help mitigate inaccuracies observed in MLLM-based assessments. Investigating the attention maps and diffusion trajectories of our model could further elucidate the underlying mechanisms of VSF. Decoupling the attention, such that it calculates the positive and negative attention separately and then uses the ratio to extrapolate the output, might yield better quality in exchange for runtime. Or, adding a scaling factor to the positive prompt for better control.
Appendix S Use of LLMs in The Paper
In this work, we used LLM for paper-related work collection and consumption. It is also used to polish the paper language or provide feedback for writing/figures. It is also used to brainstorm before and during the project.